OpenStudy (anonymous):

Find S8 for the series -2 + -10 + -50 + -250 +... i got 781248

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

it is a geometric sequence whose constant is q=5 and first term is a1=-2

OpenStudy (anonymous):

okay. and i plug it to the formula and got 781248

OpenStudy (michele_laino):

we have to apply this formula: \[\Large {S_8} = {a_1}\frac{{1 - {q^8}}}{{1 - q}} = 1 \cdot \frac{{1 - {5^8}}}{{1 - 5}} = ...?\]

OpenStudy (anonymous):

no wait it shud be 195312

OpenStudy (anonymous):

a1 should be -2 not 1

OpenStudy (michele_laino):

oops.. \[\Large {S_8} = {a_1}\frac{{1 - {q^8}}}{{1 - q}} = \left( { - 2} \right) \cdot \frac{{1 - {5^8}}}{{1 - 5}} = ...?\]

OpenStudy (anonymous):

so did i get the right anwser?

OpenStudy (michele_laino):

it is -195312 since you are summing all negative terms

OpenStudy (anonymous):

okay thank u

OpenStudy (michele_laino):

:)